OpenStudy (anonymous):

Find equation of plane that contains y-axis makes an angle of pi/6 with the positive x-axis

OpenStudy (anonymous):

I guess I could use three points. If I use (0,0,0), (1, 0,1/sqrt(3)) and (1,1, 1/sqrt(3)_

OpenStudy (amistre64):

3 points will work as long as they meet the critieria

OpenStudy (amistre64):

0,0,0 0,1,0 x,0,z where x and z are the results of the tilt

OpenStudy (anonymous):

yes, I am using x = 1 to solve for my z

OpenStudy (anonymous):

The answer I cam up with was -1/3 x - z = 0

OpenStudy (amistre64):

 my thought would amount to using 2,0,1 to avoid fractions and such

OpenStudy (amistre64):

err, sqrt(3),0,1 maybe

OpenStudy (amistre64):

forogt how to divvy up a 30 degree angle :)

OpenStudy (amistre64):

x 0 sqrt3 ; x = 1 y 1 0 ; y= -(1-sqrt3) z 0 1 ; z= -sqrt(3) x -(1-sqrt3)y -sqrt3 z = 0 but i might be rusty

OpenStudy (anonymous):